OpenStudy (anonymous):

Line m passes through the points (5, 1) and (8, 6) while line n passes through the points (-4, 3) and (-1, 8). Which statement accurately describes the relationship between the two lines? A. Lines m and n have the same slope so they are perpendicular. B. Lines m and n have the same slope so they are parallel. C. Lines m and n have opposite reciprocal slopes so they are parallel. D. Lines m and n have opposite reciprocal slopes so they are perpendicular. @soulkingdom @ZDoubleM

OpenStudy (anonymous):

i think B @ZDoubleM what do u think

OpenStudy (anonymous):

i will be back in a few have to change out of my pj

OpenStudy (anonymous):

ok @soulkingdom

OpenStudy (anonymous):

thankz

OpenStudy (anonymous):

@inkyvoyd

OpenStudy (anonymous):

@Lets_Be_Cops

OpenStudy (anonymous):

told u not to tag me anymore -.- thank you

OpenStudy (anonymous):

im back

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

sorry it tack so long

OpenStudy (anonymous):

i got the answer for it

OpenStudy (anonymous):

u did what was it

OpenStudy (anonymous):

lol i kow how to do this, but i dont feel like it. But you need to use Slope formula: slope(m) = (y2 - y1) / (x2 - x1) for both the lines first. Line 1 m = (6 - 1) / (8 - 5) m = (5) / (3) m = 5/3 Line 2 m = (8 - 3) / (-1 + 4) m = (5) / (3) m = 5/3

OpenStudy (anonymous):

They both have the same slope, figure it out from there